OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (anonymous):

@tcarroll010 @danielcox74 @wesdg1978 @GoldPhenoix @emcrazy14 @zepdrix @zackeria_x333 @damien @FriedRice @theEric

OpenStudy (ivettef365):

that is equal to y = -3x + 7 , the slope is -3 so to be perpendicular the slope needs to be opposite reciprocal, so what is the opposite reciprocal of -3

OpenStudy (anonymous):

well you know that when the line is perpendicular the slope has to be the complete opposite and if you want to get the y intercept you just plug in the points.

OpenStudy (anonymous):

-3/1?

OpenStudy (anonymous):

no because when it becomes slope intercept form it become y = -3x + 7

OpenStudy (anonymous):

so it changes into 1/3 the slope

OpenStudy (anonymous):

so the equation turns into y = 1/3x + b

OpenStudy (anonymous):

then just plug in the points and the one that you are left with is the y intercept or b

OpenStudy (anonymous):

so now what do I do?

OpenStudy (anonymous):

what is the y intercept??

OpenStudy (anonymous):

(-1)=1/3(6) + b

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

so the equation turns out to be y = 1/3x -3

OpenStudy (anonymous):

ok im still confused hold on

OpenStudy (anonymous):

why??

OpenStudy (anonymous):

and on what?

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). lets try different one I have a math test today so I wanna study really good

OpenStudy (anonymous):

ok so what do you first of all

OpenStudy (anonymous):

find the slope?

OpenStudy (anonymous):

ok what is the slope of the perpendicular equation?

OpenStudy (anonymous):

2/1?

OpenStudy (anonymous):

no you have to change the equation entirely opposite

OpenStudy (anonymous):

I mean slope

OpenStudy (anonymous):

o so 1/2?

OpenStudy (anonymous):

no negative 1/2

OpenStudy (anonymous):

I told you entirely

OpenStudy (anonymous):

opposite

OpenStudy (anonymous):

o ok I get that then what do I do

OpenStudy (anonymous):

any body home?

OpenStudy (anonymous):

hello?????

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

and then you see that you have points (6,3) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the equation is y = -1/2x + b right. you know that already, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what do you do next?

OpenStudy (anonymous):

find y?

OpenStudy (anonymous):

yes but how?

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok just plug in the coordinates (6,3) into the equation y = -1/3x

OpenStudy (anonymous):

and then the answer will be the b

OpenStudy (anonymous):

be the b?

OpenStudy (anonymous):

b is the y intercept in the equation y = mx + b

OpenStudy (anonymous):

o

OpenStudy (anonymous):

so plug in the points and tell me what you got

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

just plug in the points into the equation y = -1/3x + b and like (6,3) goes in like 3 = -1/3(6) + b

OpenStudy (anonymous):

are you still confused?

OpenStudy (anonymous):

omg I get it x and y go to x and y yes!

OpenStudy (anonymous):

ok then just do the way you will to get b

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you have the equation 3 = -1/3(6) + b right you just need to do the equation like 3 = -2 + b and then move the negative 2 to the other side then you will get b = 5

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

y = −one halfx + 6

OpenStudy (anonymous):

is this a new question?

OpenStudy (anonymous):

no one of the answers?????

OpenStudy (anonymous):

no it is y = -1/3x + 5

OpenStudy (anonymous):

that's not one of them?

OpenStudy (anonymous):

I mean y = -1/2x + 6

OpenStudy (anonymous):

yes you are right!

OpenStudy (anonymous):

we have a winner ^

OpenStudy (anonymous):

ding!!!! ding!!!! ding!!!

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

not really can you show me one more? Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).??

OpenStudy (anonymous):

buts its not that hard right just alot of work

OpenStudy (anonymous):

parallel lines have the same slope so you don't have to do so much on this question. Do you know what is the equation for this equation.

OpenStudy (anonymous):

hold on let me see if I can do that one

OpenStudy (anonymous):

3= -1/2(5) + b??

OpenStudy (anonymous):

no parallel lines have same slope but if it was perpendicular your doing it right.

OpenStudy (anonymous):

3 = 2(5) + b

OpenStudy (anonymous):

wd do I work that out?

OpenStudy (anonymous):

how*

OpenStudy (anonymous):

you know that your question said that you want a parallel equation right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so the slope has to be similar to the original slope which is 2 so you don't do anything to it

OpenStudy (anonymous):

so equation stays like y = 2x + b

OpenStudy (anonymous):

ok so now I ??

OpenStudy (anonymous):

just put in the points and you will get b or y intercept

OpenStudy (anonymous):

ok 3=2(5) + b

OpenStudy (anonymous):

yes what is b?

OpenStudy (anonymous):

a letter...just kidding

OpenStudy (anonymous):

ok so what is it?

OpenStudy (anonymous):

I don't know how to work it out is it -13?

OpenStudy (anonymous):

that's not right

OpenStudy (anonymous):

no its like this 3 = 2(5) + b right then just change it like this 3 = 10 + b then just move it like this 3 - 10 = 10 - 10 + b -7 = b

OpenStudy (anonymous):

y = 2x − 7??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you lol u have helped alot

OpenStudy (anonymous):

ok do well on your test and don't forget about parallel and perpendicular equations

OpenStudy (anonymous):

you get confused too much on those